OpenStudy (4everaddicted2anime):

Medal and Fan. A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone.

OpenStudy (retireed):

V of a cube needs to be equal to the volume of the cone. Correct?

OpenStudy (4everaddicted2anime):

I believe so

OpenStudy (retireed):

What is the formula for the volume of a cube? What is the formula for the volume of a cone?

OpenStudy (4everaddicted2anime):

Cube=Lwh Cone=pi * r^2 * l

OpenStudy (retireed):

Very good. Those will work but for this problem I like .... V = s^3 for the cube V= 1/3 r^2 h for the cone since they want to the height of the cone Set these equation equal to each other

OpenStudy (4everaddicted2anime):

2*2*2=1/3 * 3^2 * h

OpenStudy (retireed):

Yes now solve for H Vcube = s^3 = 1/3 r^2 h = Vcone

OpenStudy (4everaddicted2anime):

Okay so it is 8=3h Then I divide by 3 to get 8 divided by 3pi = h

OpenStudy (4everaddicted2anime):

Did I do it right?

OpenStudy (retireed):

I made a mistake I forgot the pi

OpenStudy (retireed):

Very good. Those will work but for this problem I like .... V = s^3 for the cube V= 1/3 pi r^2 h for the cone since they want to the height of the cone Set these equation equal to each other 2*2*2=1/3 * pi * 3^2 * h Yours was correct

OpenStudy (retireed):

8 * 3 / (pi * 9) = h

OpenStudy (4everaddicted2anime):

Okay. Thank you so much

OpenStudy (retireed):

I got h = 0.85

OpenStudy (retireed):

Sorry for my typing error

OpenStudy (4everaddicted2anime):

It's okay.